A company purchased a delivery van for $28000 With a Salvage value of $3000 on September 1 year 1 It has an estimated useful life of 5 years. Using the straight-line method how much depreciation expense should the company recognize on December 31 year 1

A) $1400 B) 2,067 C) $1667 D) $ 5000 E) $1250   

Wickland Company Installs a Manufacturing machine in its production facility at the beginning of the year at a cost of $87000 The machine Useful life is estimated to be 5 years or $400000 units of product with a $7000 salvage value. During its second year, the machine produces $84500 units of product. Determine the machines second-year depreciation under the double-declining balance method

A) $18379 B) $20880 C) $16900 D) $17400 E) $16000

A Company Purchased equipment for $65000 with a salvage value of $5000 on April 1 Year 1. It has an estimated useful life of 5 years. Using the straight-line method, how much depreciation expense should the company recognize on December 31 year 1?

A) $3000 B) $9000 C) $13000 D) $1000 E) $12000
Answer #1

1) Straightline method

annual Depreciation =(Cost -salvage ) / Life in years

Annual Depreciaion = (28000-3000)/5 = 5000

Depreicaion From Sept to December = 5000*3/12

= $ 1250 (Answer) option : E

2) Double Decline Balance method

Depreciation Rate = 1/ Life *2 = 40%

Depreciation on Second year = 87,000*(1-40%)*40%

= $20,880 (Answer) Option : B

3). Straight Line Method

Depriciation = (65000-5000)/5 =12,000

Depreciaiton for April to Dec = 12,000*9/12

= $9,000 (Answer) Option : B